# CIF-Bench: A Chinese Instruction-Following Benchmark for Evaluating the Generalizability of Large Language Models

### v1

@@ 99 Function Explanation
117 Material Synthesis
Bara et al. (2021)
118 Tool use
Taskmatrix Liang et al. (2023)
119 Concept Abstraction
N/A
120 Rhyme Aligned Generation
N/A
121 Advertising
N/A
122 Mind Tree Generation
N/A
123 First Order Logic
FOLIO Han et al. (2022)
124 Critical Thinking
N/A
125 Empathy Detection
N/A
126 Social Norms Detection
Moral Stories Emelin et al. (2021)
127 Make Positive
Ziems et al. (2022)
128 Translate to Ancient Chinese
N/A
129 Recipe Generation
Market (2018)
130 Imagination
N/A
131 Compositional Reasoning
Lake and Baroni (2017)
132 Personality Detection
N/A
133 Table Generation
N/A
134 Flowchart Generation
N/A
135 Review Generation
N/A
136 Draw Figure with symbol
N/A
137 CommonsenseQA
CommonsenseQA Talmor et al. (2018)
138 ReadingComprehensionQA
Rajpurkar et al. (2018)
139 DiscreteOperationQA
DROPDua et al. (2019)
140 MultiHopQA
HotpotQA Yang et al. (2018)
141 CommonsenseNLI
HellaSwag Zellers et al. (2019)
142 ConversationalQA
CoQA Reddy et al. (2019)
143 MathQA
GSM8K Cobbe et al. (2021)
144 English translation
N/A
145 French translation
DiaBLa Bawden et al. (2021)
146 Arabic translation
Altammami et al. (2020)
147 Japanese translation
Oda (2016)
148 Spanish translation
"European Commission and Technology." (2017)
149 Bengali translation
Islam et al. (2022)
150 Tamil translation
Ramasamy et al. (2012)
151 Gujarati translation
Shah and Bakrola (2019)
@@ Appendix B CIF-Bench Results in Public Split
We provide the category-based results in public split here.

### v2

@@ 96 Chinese Heteronomous Language Detection N/A
117 Material Synthesis Bara et al. (2021)
118 Tool use Taskmatrix Liang et al. (2023)
119 Concept Abstraction N/A
120 Rhyme Aligned Generation N/A
121 Advertising N/A
122 Mind Tree Generation N/A
123 First Order Logic FOLIO Han et al. (2022)
124 Critical Thinking N/A
125 Empathy Detection N/A
126 Social Norms Detection Moral Stories Emelin et al. (2021)
127 Make Positive Ziems et al. (2022)
128 Translate to Ancient Chinese N/A
129 Recipe Generation Market (2018)
130 Imagination N/A
131 Compositional Reasoning Lake and Baroni (2017)
132 Personality Detection N/A
133 Table Generation N/A
134 Flowchart Generation N/A
135 Review Generation N/A
136 Draw Figure with symbol N/A
137 CommonsenseQA CommonsenseQA Talmor et al. (2018)
138 ReadingComprehensionQA Rajpurkar et al. (2018)
139 DiscreteOperationQA DROPDua et al. (2019)
140 MultiHopQA HotpotQA Yang et al. (2018)
141 CommonsenseNLI HellaSwag Zellers et al. (2019)
142 ConversationalQA CoQA Reddy et al. (2019)
143 MathQA GSM8K Cobbe et al. (2021)
144 English translation N/A
145 French translation DiaBLa Bawden et al. (2021)
146 Arabic translation Altammami et al. (2020)
147 Japanese translation Oda (2016)
148 Spanish translation "European Commission and Technology." (2017)
149 Bengali translation Islam et al. (2022)
150 Tamil translation Ramasamy et al. (2012)
151 Gujarati translation Shah and Bakrola (2019)
@@ Appendix B CIF-Bench Results in Public Split
We provide the category-based results in public split here in Table 13.